Angie has 2 twin brothers, John and James. John says to her, “I am 3 times your age, Angie, plus 1 year,” and James says to her, “I am four times your age minus 3 years.” Using x to represent her age, how old is Angie? How old are her brothers? Show your work.
Answer by ankor@dixie-net.com(22740)

Angie has 2 twin brothers, John and James.
John says to her, “I am 3 times your age, Angie, plus 1 year,” and James says to her, “I am four times your age minus 3 years.”
Using x to represent her age, how old is Angie? How old are her brothers? Show your work.
:
Let b = the age of the twin brothers
let x = Angie's age
:
Write an equation for each statement
“I am 3 times your age, Angie, plus 1 year,”
b = 3x + 1
“I am four times your age minus 3 years.”
b = 4x - 3
b = b therefore
4x - 3 = 3x + 1
4x - 3x = 1 + 3
x = 4 yrs is Angie's age
:
:
you can check the solution in both statements and find the brothers' age